# Bayesian Concept Bottleneck Models with LLM Priors

### v1

@@ 5. Does the note imply the patient is retired?
CBMs with $K=6$ were trained on 100, 200, 400, and 800 observations and evaluated on 500 held-out observations. This was done for 4 replicates, where the data were randomly selected from the full dataset of 7000 observations. An Oracle CBM was also fit as the concepts are known.
@@
BC-LLM outperforms the comparator methods both in predictive performance and uncertainty quantification (Figure 2a) and, with only 400 observations, performs as well as Oracle. The performance gap between BC-LLM and the comparator methods widens with more training observations because BC-LLM can iteratively refine its concepts. Although Boosting is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations. Additionally, BC-LLM was better at recovering the true concepts compared to existing methods.
Table 1: AUCs of CBMs and black-box models (ResNet) for classifying bird subcategories.
Bird # classes BC-LLM One-pass Boosting ResNet
Albatross 3 0.837 (0.771, 0.894) 0.833 (0.771, 0.890) 0.775 (0.709, 0.842) 0.837 (0.767, 0.897)
Auklet 4 0.884 (0.830, 0.928) 0.865 (0.813, 0.914) 0.733 (0.671, 0.792) 0.831 (0.767, 0.893)
Blackbird 4 0.869 (0.822, 0.912) 0.719 (0.671, 0.771) 0.749 (0.689, 0.801) 0.857 (0.811, 0.900)
Bunting 3 0.932 (0.880, 0.972) 0.629 (0.585, 0.672) 0.681 (0.593, 0.762) 0.919 (0.869, 0.962)
Flycatcher 7 0.862 (0.835, 0.887) 0.800 (0.766, 0.833) 0.784 (0.753, 0.813) 0.817 (0.781, 0.849)
Sparrow 21 0.803 (0.785, 0.819) 0.723 (0.704, 0.739) 0.618 (0.597, 0.640) 0.812 (0.794, 0.830)
Tern 7 0.817 (0.784, 0.849) 0.705 (0.667, 0.740) 0.609 (0.567, 0.651) 0.714 (0.674, 0.751)
The posterior distribution over concepts can be visualized using hierarchical clustering (Figure 2b). Comparing concepts learned given 100 versus 800 observations, we find that the posterior learned by BC-LLM notably contracts towards the true concepts with increasing data.
4.3 Learning to classify birds
To demonstrate the multimodality of BC-LLM, we evaluate how well BC-LLM and other CBM methods can differentiate between bird categories in the CUB-birds image dataset (Welinder et al.,, 2010). [^1] We classify bird categories rather than all subtypes because there is likely a parsimonious CBM for classifying between categories but not between all subtypes. While the exact set of $K$ concepts needed to classify bird categories are unknown, we know certain features that are certainly spurious (e.g. features in the background are irrelevant).
Similar to the prior experiment, we do not tell the LLM that we are predicting bird types. By replacing the bird labels with one-hot encodings, the supervised learning task is more difficult, as the LLM cannot rely on prior knowledge about birds and must conduct an unbiased exploration over concepts. The data is split 50/50 between training and testing (so 30 bird images per category are available for training). Due to cost limitations, seven bird classification tasks were considered. The number of concepts $K$ learned for each task was set to the number of classes, but no smaller than 4 and no greater than 10. For the black-box comparator, we fine-tuned the last layer of ResNet50 pre-trained on ImageNetV2 (He et al.,, 2016; Recht et al.,, 2019).
BC-LLM outperforms all the other CBM learning procedures for every bird type (Table 1). Moreover, it outperforms ResNet50 model in all but one bird. Whereas ResNet50 easily overtrained on the small bird datasets, BC-LLM identified relevant concepts as inputs to a simple parametric model, leading to faster model convergence.
As an example, we show concepts learned by BC-LLM for classifying Bunting birds in Figure 3b. The concepts learned by BC-LLM tended to be more relevant than those from comparator methods. For instance, given all the training data, One-pass summarization checks for a “bird sitting sideways”, “bird on a perch”, and “bird feeding or foraging,” whereas BC-LLM assigned bird position a low posterior weight. Moreover, visualizing the learned concepts helps us understand the learning procedure itself. For instance, we see that BC-LLM learns to downweight spurious features, i.e. concepts relating to backgrounds or bird positions, with accumulating data.
We also assess how robust BC-LLM is in out-of-distribution (OOD) settings, as the Bayesian framing should improve uncertainty quantification. We apply the Bunting CBMs to classify images of the first author’s dog (Fig 3c). Compared to other CBM methods, BC-LLM was more appropriately unsure which Bunting bird to classify the dog as. Interestingly, when the dog is wearing a rainbow unicorn costume (C2), One-pass and Boosting overconfidently classify the dog as the colorful Bunting category (A3) with probabilities greater than 95%. Likewise, when these CBMs were used to classify images of finches (also small colorful birds), the average entropy of BC-LLM was again highest.
Method C1 C2 Finches (e.g., C3)
BC-LLM 0.91 0.90 0.70 (0.53 -0.88)
One-pass 0.06 0.20 0.27 (0.00-0.66)
Boosting 0.00 0.00 0.15 (0.00-0.57)
Figure 3: Bunting bird classification across categories A1-A3. (b) Highlighted labels in the dendrogram correspond to relevant features. Labels are shortened concept questions, generally in the format “Does the image depict…?” Numbers in parentheses describe the proportion of posterior samples with the concept. Left and right dendrograms correspond to learned posteriors for 1/3 and 3/3 of the training data, respectively. (c) To test robustness in OOD settings, the Bunting bird CBMs are asked to classify two images of the first author’s dog and 30 images of finches. Entropies of the model predictions (a measure of model uncertainty) are shown for all the CBM methods. 95% CI is shown for the Finch dataset.
Method AUC (95% CI) Log Lik (95% CI)
BC-LLM 0.64 (0.58, 0.70) -0.45 (-0.50, -0.40)
One-pass 0.59 (0.52, 0.65) -0.46 (-0.51, -0.41)
Boosting 0.59 (0.52, 0.66) -0.46 (-0.51, -0.41)
Bag-of-words 0.52 (0.46, 0.58) -0.47 (-0.52, -0.41)
Figure 4: Learning to augment a readmission risk prediction model for heart failure patients. Labels are shortened versions of concept questions of the format “Does the note mention the patient having…?”
4.4 Augmenting an existing tabular model with real-world clinical notes
@@
Finally, we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance. In particular, the data science team at the Zuckerberg San Francisco General Hospital (ZSFG) has both tabular data and unstructured clinical notes from the electronic health record (EHR). The team has so far fit a model using the tabular data to predict 30-day unplanned readmission risk for patients with congestive heart failure (CHF). Using CBMs, the team wants to assess if clinical notes contain valuable information to extract as additional features. Moreover, the learned model should be interpretable and easy to audit.
To address this, we extend BC-LLM by having the CBM take as inputs (i) the risk prediction from the original tabular model and (ii) $K$ concepts extracted from the clinical notes. This can be viewed as revising the existing model by including additional features. We set to $K=4$. The CBMs were trained on 1000 patients and evaluated on 500 held-out patients. Data in this experiment contained PHI, for which IRB approval was obtained.
The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60. BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Figure 4). Moreover, it outperforms comparator methods with respect to both AUC and log likelihood. When the CBMs were shown to four clinicians, [^2] All clinicians work or have worked in ZSFG. Three are in the hospital’s data science team, one of whom (LZ) is a coauthor on this paper. The fourth leads a separate clinical data science research group. They were all interviewed separately. they found that the CBM learned by BC-LLM to be substantially more interpretable and meaningful than the existing tabular model. The clinicians from the hospital’s data science team noted that they had not considered engineering many of the learned concepts on the list. All clinicians noted that many of the learned concepts are known to be causally relevant and would have been difficult to learn from existing tabular data. For example, “substance dependence” is rarely documented directly in the tabular data, and information relevant to determining its value is often missing. Finally, the clinicians highly valued the fact that many of the learned concepts were clinically actionable, such as helping patients better control their diabetes or opting for medication regimes that are easier for patients to follow.
@@ 5 Discussion
BC-LLM is a new method for learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations. The method is highly general: it is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification. The empirical results show that BC-LLM outperforms existing methods, learns to recover relevant concepts and downweight spurious ones, and can suggest useful features for ML developers to engineer. Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
@@ Acknowledgements
JF was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619). YT gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00. The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway. The authors thank Yifan Mai, Seth Goldman, Andrew Bishara, and Jim Marks for their invaluable feedback on the work. The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee.
@@ Appendix A Multiple-Try Metropolis-Hastings
In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3). To draw connections with the original Multiple-Try Metropolis-Hastings method (Liu et al.,, 2000), we first work in a completely abstract setting.
@@
We briefly describe Liu et al., (2000)’s original method. Suppose the current state is $x$.
@@
One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in Liu et al., (2000)). We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted. This is indeed possible.
@@
[…]
@@
If $T(-,-)$ is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
@@ Appendix B Implementation details for BC-LLM
  • Warm-start and Burn-in: Since Gibbs sampling can be slow to converge, we precede it with a warm-start, which we obtain by updating concepts greedily. That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\boldsymbol{\Phi})$, instead of sampling from the distribution. In experiments, we run warm-start for one epoch and stored the last 20 iterates as posterior samples; the rest of the samples were treated as burn-in.
@@
Appendix C Example prompt for keyphrase extraction (Step 0)
@@
Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, social determinants of health, etc). For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor. All descriptors, synonyms, and generalizations cannot be more than two words. Output as a JSON in the following format. Do not output more than 30 entries in the JSON:
Appendix D Example prompt for proposing concepts (Step 2)
The goal is to come up with a concept bottleneck model (CBM) that only extracts 5 meta-concepts from patient notes to predict some outcome Y with maximum accuracy. A meta-concept is a binary feature extractor defined by a yes/no question. We have 4 meta-concepts so far:
@@ Appendix D Example prompt for proposing concepts (Step 2)
    Does the note mention the patient having social support?
@@
    Does the note mention the patient having stable blood pressure?
  3. 3.
  4. 4.
To come up with the 5th meta-concept, I have done the following: I first fit a CBM on the 4 existing meta-concepts. Then to figure out how to improve this 4-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 4-concept CBM. These are the top extracted concepts in the resulting residual model, in descending order of importance:
$\langle$ top words from residual model $\rangle$
To interpret this residual model, a general rule of thumb is that an extracted concept with a large positive coefficient means that the outcome is positively associated with the concept being mentioned in the note, and an extracted concept with a negative coefficient means that the outcome is negatively associated with the concept being mentioned in the note. Remember that notes typically discuss the current state as well as the past medical history of a patient, so a concept being mentioned is in reference to the present OR past state of the patient.
Given the residual model, create cohesive candidates for the 5th meta-concept. Be systematic and consider all the listed concepts in the residual model. Start from the most to the least predictive concept. For each concept, check if it matches an existing meta-concept or create a new candidate meta-concept. Work down the list, iterating through each concept. Clearly state each candidate meta-concept as a yes/no question.
Suggestions for generating candidate meta-concepts: Do not propose meta-concepts that are simply a union of two different concepts (e.g. “Does the note mention this patient experiencing stomach pain or being female?” is not allowed), questions with answers that are almost always a yes (e.g. the answer to “Does the note mention this patient being sick?” is almost always yes), or questions where the yes/no options are not clearly defined (e.g. “Does the note mention this patient experiencing difficulty?” is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc). Do not propose meta-concepts where you would expect over 95% agreement or disagreement with the 4 existing meta-concepts (e.g. “Does the note mention the patient having high blood pressure?” overlaps too much with “Does the note mention the patient having hypertension?”).
Finally, summarize all the generated candidates for the 5th meta-concept in a JSON. Merge any candidate meta-concepts that are essentially the same (where you would expect over 95% agreement) or essentially opposites (you would expect over 95% disagreement). In the JSON, include a list of comma-separated list of phrases that mean the same or opposite of each candidate meta-concept.
Appendix E Example prompt for annotations for candidate concepts (Step 3)
A simple concept extraction procedure is to obtain binary or categorical concept extractions. In line with findings in Kim et al., (2023), we found that BC-LLM performed better when the LLM quantified its uncertainty when extracting concepts. So if the LLM was unsure about the concept’s value, it was allowed to return a probability. As such, we use prompts like the following:

### v2

@@ 5. Does the note imply the patient is retired?
CBMs with KK=6 were trained on 100 to 800 observations. Human(Oracle)+CBM uses true concepts.
@@
BC-LLM outperforms the comparator methods both in predictive performance and calibration (Figure 3a) and, with only 400 observations, performs as well as Human(Oracle)+CBM. The performance gap between BC-LLM and the comparator methods widens with more training observations because BC-LLM can iteratively refine its concepts. Although Boosting LLM+CBM is also an iterative procedure, it struggles to hypothesize relevant concepts because it reviews only a few misclassified observations each iteration and cannot revise concepts added in previous iterations. Additionally, BC-LLM was better at recovering the true concepts compared to existing methods.
Visualizing the posterior distribution using hierarchical clustering (Fig 3 right), the learned posterior notably contracts towards the true concepts with increasing data as the number of training observations increases from 100 to 800. Moreover, BC-LLM remains appropriately unsure about certain concepts that are highly correlated. For instance, it states many variations of the smoking concept that are essentially impossible to distinguish between, which is both expected and desired.
4.3 Augmenting a tabular model with clinical notes
@@
Here we demonstrate how BC-LLM can help ML developers integrate different data modalities to improve model performance. The data science team at hospital [Anonymous] has both tabular data and unstructured notes from the electronic health record (EHR). The team has so far fit a model using the tabular data to predict readmission risk for heart failure patients and wants to assess if clinical notes contain additional concepts that are predictive. To address this, we extend BC-LLM to take as inputs (i) the risk prediction from the original tabular model and (ii) $K=4$ concepts extracted from the clinical notes, which can be viewed as revising the existing model by adding features. The CBMs were trained on 1000 patients and evaluated on 500 held-out patients.
The original tabular model, trained on lab and flowsheet values, achieved an AUC of 0.60. BC-LLM revised this model to perform substantially better, achieving an AUC of 0.64 (Fig 4 left). Moreover, it outperforms comparator methods with respect to both AUC and brier score.
To assess the interpretability of the learned concepts, four clinicians were asked to rate concepts in the CBMs as well as top features in the original tabular model by how predictive they were, from 1==low to 3==high clinical relevance. [^4] Tabular features were rewritten as questions so they could not be distinguished from learned concepts. See Appendix for annotation instructions. Concepts from BC-LLM received an average rating of 2.5, whereas concepts from the other methods were rated $\leq$2.1 (Fig 4 right). When unblinded to the concepts, clinicians commented on a number of advantages to using BC-LLM. First, many of the learned concepts were known to be causally relevant but were difficult to engineer solely from tabular data. For example, information for determining “substance dependence” is rarely documented in the tabular data but often noted in clinical notes. Second, BC-LLM also suggested new features to engineer that the data science team had not previously considered. Most interestingly, some of the learned concepts pointed to interventions that the hospital could try to implement, such as ensuring follow-up appointments were scheduled and opting for medication regimes that were easier to follow. In contrast, the top features in the original tabular model were difficult to interpret and not actionable.
Method AUC (95% CI) Brier (95% CI)
BC-LLM 0.64 (0.58, 0.70) 0.14 (0.12, 0.62)
LLM+CBM 0.59 (0.52, 0.65) 0.29 (0.25, 0.33)
Boosting LLM+CBM 0.59 (0.52, 0.66) 0.14 (0.12, 0.16)
Bag-of-words 0.52 (0.46, 0.58) 0.29 (0.25, 0.34)
Refer to caption Refer to caption
Figure 4: Learning to augment a readmission risk prediction model for heart failure patients. Dendrogram labels are shortened questions of the format “Does the note mention the patient having…?”. Highlighted concepts received scores from clinicians as being highly predictive (scores 2.5+). Average clinician ratings for concepts/features from the different methods are shown on the right.
@@ 5 Discussion
BC-LLM is a new approach to learning CBMs that iteratively proposes concepts using an LLM within a Bayesian framework, which allows for rigorous uncertainty quantification despite LLMs being prone to error and hallucinations. Because it explores and suggests concepts in a data-adaptive manner, BC-LLM is particularly well-suited for settings with limited prior knowledge about which concepts are relevant or where the number of potentially relevant concepts is infinite. The method is compatible with various data modalities (text, images, and tabular data) and can be extended beyond the settings of binary and multiclass classification. The empirical results show that BC-LLM outperforms existing methods, even black-box models in certain settings.
Future work. BC-LLM is currently designed for learning highly interpretable CBMs where the number of concepts $K$ is typically no more than $20$. While BC-LLM can be applied to learn even more concepts, future directions can consider further speeding up posterior inference, such as through mini-batching of observations or concepts.
Impact Statement. The interpretability provided by BC-LLM additionally enables better understanding of the underlying algorithm and facilitates human oversight, which may improve the safety of AI algorithms and AI-based decision-making. The use of LLMs incurs significant computational cost and corresponding environmental impacts; however, BC-LLM is much more computationally efficient than comparable LLM-based CBMs, potentially reducing overall environmental impacts.
@@ Acknowledgements
This work was funded through a Patient-Centered Outcomes Research Institute® (PCORI®) Award (ME-2022C1-25619). The views presented in this work are solely the responsibility of the author(s) and do not necessarily represent the views of the PCORI®, its Board of Governors or Methodology Committee. The authors thank the UCSF AI Tiger Team, Academic Research Services, Research Information Technology, and the Chancellor’s Task Force for Generative AI for their technical support related to the use of Versa API gateway. The authors thank Patrick Vossler, Yifan Mai, Seth Goldman, Andrew Bishara, Jim Marks, Julian Hong, Aaron Kornblith, Nicholas Petrick, and Gene Pennello for their invaluable feedback on the work.
@@ Appendix A Multiple-Try Metropolis-Hastings
In this section, we provide intuition for the Multiple-Try Metropolis-Hastings partial posterior method (Algorithm 3). To draw connections with the original Multiple-Try Metropolis-Hastings method [45], we first work in a completely abstract setting.
@@
We briefly describe [45]’s original method. Suppose the current state is $x$.
@@
One can show that this Markov chain satisfies the detailed balance equations when marginalizing out the intermediate states $z_{1},z_{2},\ldots,z_{M},z_{1}^{*},\ldots,z_{M}$ (see Theorem 1 in [45]). We would like, however, to minimize the number of proposals necessary, so it is natural to ask whether the algorithm can be modified such that the proposals needed for the backward transitions, $z_{1}^{*},\ldots,z_{M-1}^{*}$, can be omitted. This is indeed possible.
@@
$\begin{split}\pi(x)\Phi(x,y)&=k\pi(x)T(y;x)\int\prod_{i=2}^{M}T(z_{i};x)q(y;\mathbf{z})\alpha(x,y;\mathbf{z}_{-1})d\mathbf{z}_{-1}\\ &=k\int\min\left\{\pi(y)T(x;y)\prod_{i=2}^{M}T(z_{i};y)q(x;\mathbf{z}^{*}),\pi(x)T(y;x)\prod_{i=2}^{M}T(z_{i};x)q(y;\mathbf{z})\right\}d\mathbf{z}_{-1}\\ &=\pi(y)\Phi(y,x).\qed\end{split}$ (7)
@@
If T(,)T(-,-) is invariant in the first argument (as in our application to BC-LLM), then this last formula is exactly equal to $\beta(x,y;\mathbf{z},\mathbf{z}^{*})$, the acceptance probability for the classical version, but if we were to use the same points for both the non-realized forward and backward proposals.
@@ Appendix B Implementation details for BC-LLM
  • Warm-start and Burn-in: Since Gibbs sampling can be slow to converge, we precede it with a warm-start, which we obtain by updating concepts greedily. That is, we select the concept that maximizes $\operatorname{argmax}p(\gamma|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$, instead of sampling from the distribution. In experiments, we run warm-start for one epoch and stored the last 20 iterates as posterior samples; the rest of the samples were treated as burn-in.
@@
Appendix C Example prompts
Here we provide example prompts that one may use with BC-LLM. The prompt should vary with how much prior knowledge one has about the prediction target as well as how much information one would like to reveal to the LLM about the prediction target.
C.1 Example prompt for keyphrase extraction (Step 0)
Example prompt for extracting keyphrases from images, where we reveal limited information to the LLM about the prediction target:
Given an image, your task is to brainstorm a set of descriptors that will help classify the image to its corresponding label Y. For the provided image, list as many descriptors about it. For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor. All descriptors, synonyms, and generalizations cannot be more than two words. Output characteristics that this image possess, such as "round wings" or "red head." Do not output general categories of descriptors like "wing color" or "head shape." Output at least 10 keyphrases.
Example prompt for extracting keyphrases from clinical notes, where we do not reveal information about the prediction target but we provide ideas on patient characteristics that one should extract:
@@ C.1 Example prompt for keyphrase extraction (Step 0)
Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, social determinants of health, etc). For each descriptor, also list as many descriptors that mean the same thing or generalizations of the descriptor. All descriptors, synonyms, and generalizations cannot be more than two words. Output as a JSON in the following format...
C.2 Example prompt for proposing concepts (Step 2)
Example prompt for having the LLM brainstorm candidate concepts based on top keyphrases in the keyphrase model, where we do not reveal information about the prediction target and only reveal that the dataset contains patient notes:
The goal is to come up with a concept bottleneck model (CBM) that only extracts 3 meta-concepts from patient notes to predict some outcome Y with maximum accuracy. A meta-concept is a binary feature extractor defined by a yes/no question. We have 2 meta-concepts so far:
@@ C.2 Example prompt for proposing concepts (Step 2)
To come up with the 3rd meta-concept, I have done the following: I first fit a CBM on the 2 existing meta-concepts. Then to figure out how to improve this 3-concept CBM, I first asked an LLM to extract a list of concepts that are present in each note, and then fit a linear regression model on the extracted concepts to predict the residuals of the 3-concept CBM. These are the top extracted concepts in the resulting residual model, in descending order of importance:
$\langle$ top keyphrases from keyphrase model $\rangle$
Given the residual model, create cohesive candidates for the 3rd meta-concept. Be systematic and consider all the listed concepts in the residual model. Start from the most to the least predictive concept. For each concept, check if it matches an existing meta-concept or create a new candidate meta-concept. Work down the list, iterating through each concept. Clearly state each candidate meta-concept as a yes/no question.
Suggestions for generating candidate meta-concepts: Do not propose meta-concepts that are simply a union of two different concepts (e.g. ‘‘Does the note mention this patient experiencing stomach pain or being female?" is not allowed), questions with answers that are almost always a yes (e.g. the answer to ‘‘Does the note mention this patient being sick?" is almost always yes), or questions where the yes/no options are not clearly defined (e.g. ‘‘Does the note mention this patient experiencing difficulty?" is not clearly defined because difficulty may mean financial difficulty, physical difficulties, etc). Do not propose meta-concepts where you would expect over 95% agreement or disagreement with the 4 existing meta-concepts (e.g. ‘‘Does the note mention the patient having high blood pressure?" overlaps too much with ‘‘Does the note mention the patient having hypertension?").
Finally, summarize all the generated candidate concepts in a JSON.
C.3 Example prompt for annotations for candidate concepts (Step 3)
A simple concept extraction procedure is to obtain binary or categorical concept extractions. We use prompts like this: